OpenStudy (anonymous):

Verify the trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. -tan^2x + sec^2x = 1

OpenStudy (anonymous):

@mathmath333 @mwhit2101 @Preetha @Para_phrenic88 @tinybookworm @RyGuy @waterineyes @eliassaab @goformit100 @gelow @blackops2luvr @Brandonsnider188 @Dianegrace @dancesmartgirl25 @dragonmaster917 @aaronq @Abhisar @AriPotta @Alchemista @freckles @campbell_st @Compassionate

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

\(-tan^2x+sec^2x=1\) \(\Longleftrightarrow ~sec^2x=1+tan^2x\) \(\Longleftrightarrow ~ \Large \frac {1}{cos^2x} \normalsize= 1+ \Large \frac{sin^2s}{cos^2x}\) \(\Longleftrightarrow ~ \Large \frac {1}{cos^2x} \normalsize=\Large \frac{cos^2x+sin^2s}{cos^2x}\) (true)

OpenStudy (campbell_st):

well use - tan^2 = - sin^2/cos^2 and sec^2 = 1/cos^2

OpenStudy (campbell_st):

you have the common denominator... of cos^2 so you have (-sin^2 + 1)/cos^2

OpenStudy (campbell_st):

rewriting the numerator and you get \[\frac{1 - \sin^2(x)}{\cos^2(x)}\] use a well know trig identity and substitute it into the numerator...

OpenStudy (anonymous):

I understand everything but the last part. What do you mean a well known trig identity?

OpenStudy (campbell_st):

well the best known trig identity is \[\sin^2(x) + \cos^2(x) = 1\] just make cos^2 the subject and it should be obvious

OpenStudy (anonymous):

I'm stuck at (-sin^2x + 1)/cos^2x = 1

OpenStudy (campbell_st):

dang \[\sin^2(x) + \cos^2(x) = 1 ~~~~~then~~~~~\cos^2(x) = -\sin(x) + 1\]

OpenStudy (campbell_st):

substitute this into the numerator...

OpenStudy (anonymous):

you mean denominator?

OpenStudy (anonymous):

ohh

OpenStudy (campbell_st):

either substitution works... I'd substitute the numerator

OpenStudy (anonymous):

I got it. Thank you so much.

OpenStudy (campbell_st):

and for my 3rd try its\[\sin^2(x) + \cos^2(x) = 1~~~then~~~~\cos^2(x) = -\sin^2(x) + 1\]